# Learning Performance Maximizing Ensembles with Explainability Guarantees

### v2

@@ Results
Figure 2: This figure shows two examples of the explainability vs performance trade-off, comparing the random, oracle, and learned allocation curves. The PolR dataset is an example of complementary $g$ and $b$ models, resulting in an allocated ensemble that outperforms both component models across most of the $q$ range. The SuperconductR dataset is an example of an explainability “free lunch” in which the $b$ accuracy is maintained while increasing explainability using the allocator. Curves for all datasets are available in the Appendix of the long form paper available on arxiv.org.
Dataset
AUC
PPCR
PQEOM
PQOM
PCFA
95TQM
Max Acc
Argmax $q$
$s$ Acc
Wine
79 $\pm$ 0
21 $\pm$ 0
71 $\pm$ 0
0 $\pm$ 0
7 $\pm$ 0
98 $\pm$ 0
80 $\pm$ 0
70 $\pm$ 0
78 $\pm$ 0
Phoneme
87 $\pm$ 0
12 $\pm$ 1
78 $\pm$ 4
7 $\pm$ 2
6 $\pm$ 1
100 $\pm$ 0
87 $\pm$ 0
50 $\pm$ 34
81 $\pm$ 2
KDDIPUMS
88 $\pm$ 0
17 $\pm$ 1
65 $\pm$ 3
34 $\pm$ 5
17 $\pm$ 7
100 $\pm$ 0
88 $\pm$ 0
66 $\pm$ 9
80 $\pm$ 1
EyeMovements
66 $\pm$ 0
33 $\pm$ 0
55 $\pm$ 1
7 $\pm$ 6
15 $\pm$ 2
70 $\pm$ 0
68 $\pm$ 0
31 $\pm$ 24
52 $\pm$ 0
Pol
98 $\pm$ 0
49 $\pm$ 0
98 $\pm$ 0
2 $\pm$ 0
0 $\pm$ 0
100 $\pm$ 0
99 $\pm$ 0
98 $\pm$ 0
96 $\pm$ 0
Bank
76 $\pm$ 0
-19 $\pm$ 0
4 $\pm$ 1
0 $\pm$ 0
0 $\pm$ 0
100 $\pm$ 0
79 $\pm$ 0
100 $\pm$ 0
71 $\pm$ 0
MagicTelescope
86 $\pm$ 0
39 $\pm$ 0
87 $\pm$ 2
12 $\pm$ 11
10 $\pm$ 0
100 $\pm$ 0
86 $\pm$ 0
47 $\pm$ 28
82 $\pm$ 1
House16H
89 $\pm$ 0
40 $\pm$ 0
84 $\pm$ 4
9 $\pm$ 6
6 $\pm$ 2
98 $\pm$ 0
89 $\pm$ 0
82 $\pm$ 8
86 $\pm$ 0
Credit
78 $\pm$ 0
5 $\pm$ 1
56 $\pm$ 4
14 $\pm$ 19
95 $\pm$ 0
100 $\pm$ 0
78 $\pm$ 0
76 $\pm$ 6
72 $\pm$ 0
California
90 $\pm$ 0
52 $\pm$ 0
88 $\pm$ 0
0 $\pm$ 0
7 $\pm$ 0
98 $\pm$ 0
91 $\pm$ 0
88 $\pm$ 0
90 $\pm$ 0
Electricity
92 $\pm$ 0
58 $\pm$ 0
88 $\pm$ 0
0 $\pm$ 0
7 $\pm$ 0
98 $\pm$ 0
93 $\pm$ 0
88 $\pm$ 0
92 $\pm$ 0
Jannis
79 $\pm$ 0
30 $\pm$ 0
53 $\pm$ 5
21 $\pm$ 5
14 $\pm$ 1
98 $\pm$ 0
79 $\pm$ 0
35 $\pm$ 6
76 $\pm$ 0
MiniBooNE
94 $\pm$ 0
54 $\pm$ 0
90 $\pm$ 0
0 $\pm$ 0
5 $\pm$ 0
100 $\pm$ 0
94 $\pm$ 0
90 $\pm$ 0
92 $\pm$ 0
WineR
73 $\pm$ 0
43 $\pm$ 0
85 $\pm$ 0
10 $\pm$ 0
20 $\pm$ 0
90 $\pm$ 0
74 $\pm$ 0
78 $\pm$ 0
67 $\pm$ 0
IsoletR
91 $\pm$ 0
89 $\pm$ 0
68 $\pm$ 0
0 $\pm$ 0
49 $\pm$ 10
73 $\pm$ 0
95 $\pm$ 0
68 $\pm$ 0
87 $\pm$ 1
CPUR
75 $\pm$ 0
40 $\pm$ 2
52 $\pm$ 18
0 $\pm$ 0
29 $\pm$ 11
80 $\pm$ 0
77 $\pm$ 0
70 $\pm$ 0
59 $\pm$ 1
SulfurR
98 $\pm$ 0
63 $\pm$ 2
73 $\pm$ 9
1 $\pm$ 2
79 $\pm$ 4
100 $\pm$ 0
98 $\pm$ 0
84 $\pm$ 22
96 $\pm$ 0
BrazilianHousesR
96 $\pm$ 0
71 $\pm$ 1
83 $\pm$ 0
64 $\pm$ 6
1 $\pm$ 2
93 $\pm$ 0
98 $\pm$ 0
8 $\pm$ 3
88 $\pm$ 0
AileronsR
75 $\pm$ 0
13 $\pm$ 0
70 $\pm$ 10
5 $\pm$ 7
4 $\pm$ 2
100 $\pm$ 0
76 $\pm$ 0
40 $\pm$ 35
64 $\pm$ 0
MiamiHousingR
76 $\pm$ 0
44 $\pm$ 0
76 $\pm$ 0
0 $\pm$ 0
67 $\pm$ 10
80 $\pm$ 0
78 $\pm$ 0
75 $\pm$ 0
65 $\pm$ 0
PolR
88 $\pm$ 0
44 $\pm$ 1
96 $\pm$ 1
93 $\pm$ 1
88 $\pm$ 0
100 $\pm$ 0
88 $\pm$ 0
81 $\pm$ 2
84 $\pm$ 0
ElevatorsR
75 $\pm$ 0
25 $\pm$ 0
64 $\pm$ 4
2 $\pm$ 2
17 $\pm$ 0
88 $\pm$ 0
75 $\pm$ 0
51 $\pm$ 29
63 $\pm$ 0
BikeSharingR
77 $\pm$ 0
26 $\pm$ 1
88 $\pm$ 4
22 $\pm$ 16
82 $\pm$ 1
100 $\pm$ 0
78 $\pm$ 0
21 $\pm$ 13
72 $\pm$ 0
FifaR
77 $\pm$ 0
33 $\pm$ 0
95 $\pm$ 0
93 $\pm$ 0
78 $\pm$ 0
100 $\pm$ 0
77 $\pm$ 0
50 $\pm$ 0
69 $\pm$ 0
CaliforniaR
78 $\pm$ 0
38 $\pm$ 0
73 $\pm$ 0
68 $\pm$ 0
68 $\pm$ 0
93 $\pm$ 0
79 $\pm$ 0
42 $\pm$ 0
63 $\pm$ 0
HousesR
78 $\pm$ 0
41 $\pm$ 0
73 $\pm$ 0
0 $\pm$ 0
48 $\pm$ 1
80 $\pm$ 0
79 $\pm$ 0
73 $\pm$ 0
62 $\pm$ 0
SuperconductR
83 $\pm$ 0
42 $\pm$ 0
60 $\pm$ 13
24 $\pm$ 11
95 $\pm$ 0
95 $\pm$ 0
83 $\pm$ 0
57 $\pm$ 1
76 $\pm$ 0
HouseSalesR
76 $\pm$ 0
50 $\pm$ 1
64 $\pm$ 12
0 $\pm$ 0
56 $\pm$ 6
85 $\pm$ 0
78 $\pm$ 0
78 $\pm$ 0
63 $\pm$ 0
House16HR
92 $\pm$ 0
55 $\pm$ 0
90 $\pm$ 0
0 $\pm$ 0
6 $\pm$ 0
98 $\pm$ 0
92 $\pm$ 0
90 $\pm$ 0
86 $\pm$ 0
DiamondsR
70 $\pm$ 0
19 $\pm$ 0
83 $\pm$ 0
34 $\pm$ 6
73 $\pm$ 0
100 $\pm$ 0
71 $\pm$ 0
20 $\pm$ 6
65 $\pm$ 0
MedicalChargesR
86 $\pm$ 0
24 $\pm$ 0
91 $\pm$ 1
85 $\pm$ 1
0 $\pm$ 0
100 $\pm$ 0
86 $\pm$ 0
67 $\pm$ 2
83 $\pm$ 0
@@
Thus, the following study was conducted to determine whether a consistently best feature set exists for the tabular data context used in the experiments. First, the universe of candidate features was selected to be the original features $x$ used as inputs for the ensemble component models, the predictions of both of these models $g(x)$ and $b(x)$, and finally two measures of discrepancy between the predictions, the cross-entropy $d_{ce}(g(x),b(x))$ and the mean squared error $d_{mse}(g(x),b(x))$. The measures of disagreement were included as features as they translate to the “feature independent” strategy of allocation to model $b$ for low values of $d(a(x),b(x))$ in other words the optimal allocation strategy assuming $a$ is always correct.
@@
The results support the following three conclusions. First, no one feature set proved universally best across the tested datasets and thus a full search across feature sets would be advised in settings without resource constraint. Second, although no universally best feature set was found, the “kitchen sink” set of all candidate features ($x$, $g$, $b$, $d_{ce}$, $d_{mse}$) was found to be best most consistently and was thus used to train all allocators reported in Table 1. Finally, allocators trained on just the original features $x$ were found to be consistently worst among all alternatives thus supporting the augmentation of the original features in some form. This finding is consistent with the intuition that the predictions of the component models would be very useful to learning the optimal allocation and would be either very difficult or impossible to learn from $r^{\prime}$, the optimal allocation ranking response, alone during training.
Feature Set
$\alpha:0.01$ $\alpha:0.05$ $\alpha:0.1$
$x$
18.75% 18.75% 12.50%
$g$, $b$
43.75% 43.75% 31.25%
$d_{ce}$
31.25% 18.75% 18.75%
$d_{mse}$
50.00% 43.75% 37.50%
$x$, $d_{ce}$
37.50% 25.00% 18.75%
$x$, $d_{mse}$
56.25% 37.50% 25.00%
$g$, $b$, $d_{ce}$
31.25% 25.00% 25.00%
$g$, $b$, $d_{mse}$
50.00% 50.00% 37.50%
$x$, $g$, $b$
56.25% 43.75% 37.50%
$x$, $g$, $b$, $d_{ce}$
50.00% 43.75% 37.50%
$x$, $g$, $b$, $d_{mse}$
56.25% 37.50% 37.50%
$x$, $g$, $b$, $d_{ce}$, $d_{mse}$
75.00% 56.25% 43.75%
@@
Dataset
$g_{I}$
$b_{I}$
$g_{C}$
$b_{C}$
Match?
AUC $\Delta$
Wine
Tree
DNN
Tree
DNN
Yes
0.0
Phoneme
Tree
GBT
Tree
DNN
No
KDDIPUMS
Tree
GBT
Tree
GBT
Yes
0.0
EyeMovements
Regr.
GBT
Regr.
GBT
Yes
0.0
Pol
Tree
DNN
Tree
GBT
No
Bank
Tree
DNN
Tree
DNN
Yes
0.0
MagicTelescope
Tree
GBT
Tree
GBT
Yes
0.0
House16H
Tree
GBT
Tree
GBT
Yes
0.0
Credit
Tree
GBT
Tree
GBT
Yes
0.0
California
Tree
GBT
Tree
GBT
Yes
0.0
Electricity
Tree
GBT
Tree
GBT
Yes
0.0
Jannis
Tree
GBT
Tree
GBT
Yes
0.0
MiniBooNE
Tree
GBT
Tree
GBT
Yes
0.0
WineR
Regr.
GBT
Regr.
GBT
Yes
0.0
IsoletR
Regr.
DNN
Tree
DNN
No
0.15
CPUR
Tree
GBT
Tree
DNN
No
0.0
SulfurR
Tree
DNN
Tree
GBT
No
0.04
BrazilianHousesR
Tree
GBT
Tree
DNN
No
0.08
AileronsR
Regr.
GBT
Regr.
DNN
No
0.0
MiamiHousingR
Tree
GBT
Tree
GBT
Yes
0.0
PolR
Tree
DNN
Tree
GBT
No
ElevatorsR
Regr.
DNN
Regr.
GBT
No
0.01
BikeSharingR
Tree
GBT
Tree
GBT
Yes
0.0
FifaR
Tree
GBT
Tree
DNN
No
0.0
CaliforniaR
Tree
GBT
Tree
DNN
No
0.01
HousesR
Tree
GBT
Tree
DNN
No
0.01
SuperconductR
Tree
GBT
Tree
GBT
Yes
0.0
HouseSalesR
Tree
GBT
Tree
GBT
Yes
0.0
House16HR
Tree
GBT
Tree
GBT
Yes
0.0
DiamondsR
Tree
GBT
Tree
GBT
Yes
0.0
MedicalChargesR
Tree
GBT
Tree
DNN
No
0.0
Median/Average Tree (83.9%) GBT (77.4%) Tree (87.1%) GBT (64.5%) Yes (58.1%) 0.01 $\pm$ 0.03
Table 3: For each dataset, this table reports the glass box ($g$) and black box ($b$) models which have the highest individual validation performance (Indiv. $g$ and Indiv. $b$) as well as the ($g$, $b$) pair that produces the highest allocator validation set performance (Combi. $g$ and Combi. $b$). Also shown is the difference ($AUC_{Indiv.}-AUC_{Combi.}$) in the test performance between allocators trained on the individually selected component models and paired component models - a positive value indicating that the paired component models result in a better allocator.
Acknowledgements / Thanks
@@ Appendix A Appendix
$Z^{\prime}_{\bigtriangleup\square}=Z^{\prime}_{\bigtriangleup}\cup Z^{\prime}_{\square}$ and $Z^{\prime}_{\bigtriangleup,\square}=Z^{\prime}_{\bigtriangleup}\times Z^{\prime}_{\square}$
@@
(C Sets)
$C_{\bigtriangleup,\square}(a)=\{(z_{i},z_{j}):z_{i}\in Z^{\prime}_{\bigtriangleup},z_{j}\in Z^{\prime}_{\square},a(z_{i})<a(z_{j})\}$
@@
(Increasing in $\tilde{r}$) $\forall(z_{b},z_{0},z_{2},z_{g})\in(Z^{\prime}_{b}\times Z^{\prime}_{0}\times Z^{\prime}_{2}\times Z^{\prime}_{g}),\tilde{r}^{\prime}(z_{b})<\tilde{r}^{\prime}(z_{0})<\tilde{r}^{\prime}(z_{2})<\tilde{r}^{\prime}(z_{g})$
@@
Notice $\{\tilde{r}^{\prime}(z_{b})\colon z_{b}\in Z^{\prime}_{b}\}\subseteq(-2,-1),\{\tilde{r}^{\prime}(z_{0})\colon z_{0}\in Z^{\prime}_{0}\}\subseteq(-1,0),\{\tilde{r}^{\prime}(z_{2})\colon z_{2}\in Z^{\prime}_{2}\}\subseteq(0,1),\{\tilde{r}^{\prime}(z_{g})\colon z_{g}\in Z^{\prime}_{g}\}\subseteq(1,2)$, statement follows. ∎
@@
($a^{\prime}_{q}$ to $\tilde{r}$) $|Z^{\prime}_{q}|=n_{q},Z^{\prime}_{q}=\tilde{R}^{\prime}_{q}$ where $Z^{\prime}_{q}=\{z\colon a^{\prime}_{q}(z)=1\},\tilde{R}^{\prime}_{q}=\{z\colon\mathrm{rank}(\tilde{r}^{\prime}(z))>n_{1-q}\}$
@@
Part 1: $Z^{\prime}_{q}=\{z\colon a^{\prime}_{q}(z)=1\}=\{z\colon r^{\prime}(z)>1-q\}=\{z\colon\frac{\mathrm{rank}(\tilde{r}^{\prime}(z))}{n}>1-q\}=\{z:\mathrm{rank}(\tilde{r}^{\prime}(z))>n(1-q)\}=\{z:\mathrm{rank}(\tilde{r}^{\prime}(z))>n_{1-q}\}=\tilde{R}^{\prime}_{q}$
Part 2: Observe that, by construction, $\tilde{r}^{\prime}(z)$ is the $\tilde{r}^{\prime}(z)^{th}$ percentile. Thus, $|Z^{\prime}_{q}|=|\{z\colon r^{\prime}(z)>1-q\}|=n(1-(1-q))=n_{q}$. ∎
Lemma 3.
(Set Equivalence 1) $A^{*}_{q}=B^{*}_{q},\forall q\in\mathcal{Q}$, where $A^{*}_{q}=\{a^{*}_{q}\colon a^{*}_{q}=\arg\max_{a_{q}\in A_{q}}\frac{1}{n}\sum_{i=1}^{n}s^{\prime}_{g}(z_{i})a_{q}(z_{i})+s^{\prime}_{b}(z_{i})(1-a_{q}(z_{i}))\}$ and $B^{*}_{q}=\{b^{*}_{q}:b^{*}_{q}\in A_{q},C_{g,02}(b^{*}_{q})=C_{02,b}(b^{*}_{q})=C_{g,b}(b^{*}_{q})=\emptyset\}$

### v3

@@ Results
Figure 3: This figure shows two examples of the explainability (x-axis) vs sufficient performance (y-axis) trade-off, comparing the random (blue), oracle (orange), and learned (green) allocation curves. The PolR dataset is an example of complementary $g$ and $b$ models, resulting in an allocated ensemble that outperforms both component models across most of the $q$ range. The SuperconductR dataset is an example of an explainability “free lunch” in which the $b$ accuracy is maintained while increasing explainability using the allocator. Curves for all datasets are available in Section D of the Appendix
Dataset AUC PPCR PQEOM PQOM PCFA 95TQM Max Acc Argmax $q$ $s$ Acc
Wine 79 $\pm$ 0 21 $\pm$ 0 71 $\pm$ 0 0 $\pm$ 0 7 $\pm$ 0 98 $\pm$ 0 80 $\pm$ 0 70 $\pm$ 0 78 $\pm$ 0
Phoneme 87 $\pm$ 0 12 $\pm$ 1 78 $\pm$ 4 7 $\pm$ 2 6 $\pm$ 1 100 $\pm$ 0 87 $\pm$ 0 50 $\pm$ 34 81 $\pm$ 2
KDDIPUMS 88 $\pm$ 0 17 $\pm$ 1 65 $\pm$ 3 34 $\pm$ 5 17 $\pm$ 7 100 $\pm$ 0 88 $\pm$ 0 66 $\pm$ 9 80 $\pm$ 1
EyeMovements 66 $\pm$ 0 33 $\pm$ 0 55 $\pm$ 1 7 $\pm$ 6 15 $\pm$ 2 70 $\pm$ 0 68 $\pm$ 0 31 $\pm$ 24 52 $\pm$ 0
Pol 98 $\pm$ 0 49 $\pm$ 0 98 $\pm$ 0 2 $\pm$ 0 0 $\pm$ 0 100 $\pm$ 0 99 $\pm$ 0 98 $\pm$ 0 96 $\pm$ 0
Bank 76 $\pm$ 0 -19 $\pm$ 0 4 $\pm$ 1 0 $\pm$ 0 0 $\pm$ 0 100 $\pm$ 0 79 $\pm$ 0 100 $\pm$ 0 71 $\pm$ 0
MagicTelescope 86 $\pm$ 0 39 $\pm$ 0 87 $\pm$ 2 12 $\pm$ 11 10 $\pm$ 0 100 $\pm$ 0 86 $\pm$ 0 47 $\pm$ 28 82 $\pm$ 1
House16H 89 $\pm$ 0 40 $\pm$ 0 84 $\pm$ 4 9 $\pm$ 6 6 $\pm$ 2 98 $\pm$ 0 89 $\pm$ 0 82 $\pm$ 8 86 $\pm$ 0
Credit 78 $\pm$ 0 5 $\pm$ 1 56 $\pm$ 4 14 $\pm$ 19 95 $\pm$ 0 100 $\pm$ 0 78 $\pm$ 0 76 $\pm$ 6 72 $\pm$ 0
California 90 $\pm$ 0 52 $\pm$ 0 88 $\pm$ 0 0 $\pm$ 0 7 $\pm$ 0 98 $\pm$ 0 91 $\pm$ 0 88 $\pm$ 0 90 $\pm$ 0
Electricity 92 $\pm$ 0 58 $\pm$ 0 88 $\pm$ 0 0 $\pm$ 0 7 $\pm$ 0 98 $\pm$ 0 93 $\pm$ 0 88 $\pm$ 0 92 $\pm$ 0
Jannis 79 $\pm$ 0 30 $\pm$ 0 53 $\pm$ 5 21 $\pm$ 5 14 $\pm$ 1 98 $\pm$ 0 79 $\pm$ 0 35 $\pm$ 6 76 $\pm$ 0
MiniBooNE 94 $\pm$ 0 54 $\pm$ 0 90 $\pm$ 0 0 $\pm$ 0 5 $\pm$ 0 100 $\pm$ 0 94 $\pm$ 0 90 $\pm$ 0 92 $\pm$ 0
WineR 73 $\pm$ 0 43 $\pm$ 0 85 $\pm$ 0 10 $\pm$ 0 20 $\pm$ 0 90 $\pm$ 0 74 $\pm$ 0 78 $\pm$ 0 67 $\pm$ 0
IsoletR 91 $\pm$ 0 89 $\pm$ 0 68 $\pm$ 0 0 $\pm$ 0 49 $\pm$ 10 73 $\pm$ 0 95 $\pm$ 0 68 $\pm$ 0 87 $\pm$ 1
CPUR 75 $\pm$ 0 40 $\pm$ 2 52 $\pm$ 18 0 $\pm$ 0 29 $\pm$ 11 80 $\pm$ 0 77 $\pm$ 0 70 $\pm$ 0 59 $\pm$ 1
SulfurR 98 $\pm$ 0 63 $\pm$ 2 73 $\pm$ 9 1 $\pm$ 2 79 $\pm$ 4 100 $\pm$ 0 98 $\pm$ 0 84 $\pm$ 22 96 $\pm$ 0
BrazilianHousesR 96 $\pm$ 0 71 $\pm$ 1 83 $\pm$ 0 64 $\pm$ 6 1 $\pm$ 2 93 $\pm$ 0 98 $\pm$ 0 8 $\pm$ 3 88 $\pm$ 0
AileronsR 75 $\pm$ 0 13 $\pm$ 0 70 $\pm$ 10 5 $\pm$ 7 4 $\pm$ 2 100 $\pm$ 0 76 $\pm$ 0 40 $\pm$ 35 64 $\pm$ 0
MiamiHousingR 76 $\pm$ 0 44 $\pm$ 0 76 $\pm$ 0 0 $\pm$ 0 67 $\pm$ 10 80 $\pm$ 0 78 $\pm$ 0 75 $\pm$ 0 65 $\pm$ 0
PolR 88 $\pm$ 0 44 $\pm$ 1 96 $\pm$ 1 93 $\pm$ 1 88 $\pm$ 0 100 $\pm$ 0 88 $\pm$ 0 81 $\pm$ 2 84 $\pm$ 0
ElevatorsR 75 $\pm$ 0 25 $\pm$ 0 64 $\pm$ 4 2 $\pm$ 2 17 $\pm$ 0 88 $\pm$ 0 75 $\pm$ 0 51 $\pm$ 29 63 $\pm$ 0
BikeSharingR 77 $\pm$ 0 26 $\pm$ 1 88 $\pm$ 4 22 $\pm$ 16 82 $\pm$ 1 100 $\pm$ 0 78 $\pm$ 0 21 $\pm$ 13 72 $\pm$ 0
FifaR 77 $\pm$ 0 33 $\pm$ 0 95 $\pm$ 0 93 $\pm$ 0 78 $\pm$ 0 100 $\pm$ 0 77 $\pm$ 0 50 $\pm$ 0 69 $\pm$ 0
CaliforniaR 78 $\pm$ 0 38 $\pm$ 0 73 $\pm$ 0 68 $\pm$ 0 68 $\pm$ 0 93 $\pm$ 0 79 $\pm$ 0 42 $\pm$ 0 63 $\pm$ 0
HousesR 78 $\pm$ 0 41 $\pm$ 0 73 $\pm$ 0 0 $\pm$ 0 48 $\pm$ 1 80 $\pm$ 0 79 $\pm$ 0 73 $\pm$ 0 62 $\pm$ 0
SuperconductR 83 $\pm$ 0 42 $\pm$ 0 60 $\pm$ 13 24 $\pm$ 11 95 $\pm$ 0 95 $\pm$ 0 83 $\pm$ 0 57 $\pm$ 1 76 $\pm$ 0
HouseSalesR 76 $\pm$ 0 50 $\pm$ 1 64 $\pm$ 12 0 $\pm$ 0 56 $\pm$ 6 85 $\pm$ 0 78 $\pm$ 0 78 $\pm$ 0 63 $\pm$ 0
House16HR 92 $\pm$ 0 55 $\pm$ 0 90 $\pm$ 0 0 $\pm$ 0 6 $\pm$ 0 98 $\pm$ 0 92 $\pm$ 0 90 $\pm$ 0 86 $\pm$ 0
DiamondsR 70 $\pm$ 0 19 $\pm$ 0 83 $\pm$ 0 34 $\pm$ 6 73 $\pm$ 0 100 $\pm$ 0 71 $\pm$ 0 20 $\pm$ 6 65 $\pm$ 0
MedicalChargesR 86 $\pm$ 0 24 $\pm$ 0 91 $\pm$ 1 85 $\pm$ 1 0 $\pm$ 0 100 $\pm$ 0 86 $\pm$ 0 67 $\pm$ 2 83 $\pm$ 0
@@
Thus, the following study was conducted to determine whether a consistently best feature set exists for the tabular data context used in the experiments. First, the universe of candidate features was selected to be the original features $x$ used as inputs for the ensemble component models, the predictions of both of these models $g(x)$ and $b(x)$, and finally two measures of discrepancy between the predictions, the cross-entropy $d_{ce}(g(x),b(x))$ and the mean squared error $d_{mse}(g(x),b(x))$. The measures of disagreement were included as features as they translate to the “feature independent” strategy of allocation to model $b$ for low values of $d(a(x),b(x))$ - in other words the optimal allocation strategy assuming $a$ is always correct.
@@
The results support the following three conclusions. First, no one feature set proved universally best across the tested datasets and thus a full search across feature sets would be advised in settings without resource constraint. Second, although no universally best feature set was found, the “kitchen sink” set of all candidate features ($x$, $g$, $b$, $d_{ce}$, $d_{mse}$) was found to be best most consistently and was thus used to train all allocators reported in Table 2. Finally, allocators trained on just the original features $x$ were found to be consistently worst among all alternatives thus supporting the augmentation of the original features in some form. This finding is consistent with the intuition that the predictions of the component models would be very useful to learning the optimal allocation and would be either very difficult or impossible to learn from $r$, the optimal allocation ranking response, alone during training.
Feature Set $\alpha:0.01$ $\alpha:0.05$ $\alpha:0.1$
$x$ 18.75% 18.75% 12.50%
$g$, $b$ 43.75% 43.75% 31.25%
$d_{ce}$ 31.25% 18.75% 18.75%
$d_{mse}$ 50.00% 43.75% 37.50%
$x$, $d_{ce}$ 37.50% 25.00% 18.75%
$x$, $d_{mse}$ 56.25% 37.50% 25.00%
$g$, $b$, $d_{ce}$ 31.25% 25.00% 25.00%
$g$, $b$, $d_{mse}$ 50.00% 50.00% 37.50%
$x$, $g$, $b$ 56.25% 43.75% 37.50%
$x$, $g$, $b$, $d_{ce}$ 50.00% 43.75% 37.50%
$x$, $g$, $b$, $d_{mse}$ 56.25% 37.50% 37.50%
$x$, $g$, $b$, $d_{ce}$, $d_{mse}$ 75.00% 56.25% 43.75%
@@
Acknowledgements
@@ A. Theoretical Results
$Z_{0}=\{z:z\in\mathcal{Z},s_{g}(z)+s_{b}(z)=0\}$
$Z_{2}=\{z:z\in\mathcal{Z},s_{g}(z)+s_{b}(z)=2\}$
$Z_{g}=\{z:z\in\mathcal{Z},s_{g}(z)=1,s_{b}(z)=0\}$
$Z_{b}=\{z:z\in\mathcal{Z},s_{g}(z)=0,s_{b}(z)=1\}$
$Z_{\bigtriangleup,\square}=Z_{\bigtriangleup}\times Z_{\square}$
@@
$S(a)=\frac{1}{n}\sum^{n}_{i=1}s(a,v_{i})$
$s(a,v)=s_{g}(z)a(v)+s_{b}(z)(1-a(v))$
$S_{g}(a)=\frac{1}{n}\sum^{n}_{i=1}s_{g}(z_{i})a(v_{i})$
@@
($Z$ Sets Ordered By Ranking)
$\forall(z_{b},z_{0},z_{2},z_{g})\in(Z_{b}\times Z_{0}\times Z_{2}\times Z_{g}),r(z_{b})<r(z_{0})<r(z_{2})<r(z_{g})$
@@
Notice $\{r(z_{b})\colon z_{b}\in Z_{b}\}\subseteq(-2,-1],\{r(z_{0})\colon z_{0}\in Z_{0}\}\subseteq(-1,0],\{r(z_{2})\colon z_{2}\in Z_{2}\}\subseteq(0,1],\{r(z_{g})\colon z_{g}\in Z_{g}\}\subseteq(1,2]$, statement follows. ∎
@@
(Allocated By Top Ranked)
$|V^{\prime}_{q}|=n_{q},V^{\prime}_{q}=R_{q}$ where $V^{\prime}_{q}=\{v\colon a^{\prime}_{q}(v)=1\},R_{q}=\{z\colon\mathrm{rank}(r(z))>n_{1-q}\}$
@@
Part 1: $V^{\prime}_{q}=\{v\colon a^{\prime}_{q}(v)=1\}=\{z\colon\tilde{r}(z)>1-q\}=\{z\colon\frac{\mathrm{rank}(r(z))}{n}>1-q\}=\{z:\mathrm{rank}(r(z))>n(1-q)\}=\{z:\mathrm{rank}(r(z))>n_{1-q}\}=R_{q}$
Part 2: Observe that, by construction, $\tilde{r}(z)$ is the $\tilde{r}(z)^{th}$ percentile. Thus, $|V^{\prime}_{q}|=|\{z\colon\tilde{r}(z)>1-q\}|=n(1-(1-q))=n_{q}$. ∎
Proposition 1.
(Maximal Sufficient Performance)
$\forall q\in\mathcal{Q},a^{\prime}_{q}\in A^{*}_{q}$ where $A^{*}_{q}=\{a^{*}_{q}\colon a^{*}_{q}=\arg\max_{a_{q}\in A_{q}}S(a_{q})\}$